# ThinkEdit: Interpretable Weight Editing to Mitigate Overly Short Thinking in Reasoning Models

### v3

@@ 1 Introduction
Recently, Reinforcement Learning (RL) has been applied to enhance Large Language Models (LLMs), equipping them with strong chain-of-thought (CoT) reasoning abilities Guo et al. (2025). These models, often referred to as reasoning models, first generate an intermediate reasoning process—a "thinking step"—where they reason step-by-step and then self-correct before producing a final response. As a result, they achieve remarkable improvement on mathematical reasoning tasks and demonstrate a strong ability to generate detailed CoT reasoning Jaech et al. (2024); Guo et al. (2025); Muennighoff et al. (2025).
@@ 3.2 Extracting Reasoning Length Directions
$\overline{r}^{x}_{\ell,y}=\frac{1}{|\mathcal{D}_{y}|}\sum_{i\in\mathcal{D}_{y}}\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}r^{x}_{\ell}(i,t),\vspace{-5pt}$
@@
These two vectors, $v^{\text{attn}}_{\ell}$ and $v^{\text{mlp}}_{\ell}$, capture how the model’s representation differs when reasoning chains are notably longer or shorter. In the next section, we analyze how modifying these directions in the residual stream influences both reasoning length and overall model performance.
@@ 3.3 Effects of Reasoning-Length Direction
Let $\alpha$ be a scalar weight in the range [0.08, 0.08]0.080.08[-0.08,\,0.08][ - 0.08 , 0.08 ]. For each layer $\ell$, we apply the following transformations:
$r^{\text{attn}}_{\ell}\leftarrow r^{\text{attn}}_{\ell}+\alpha\,v^{\text{attn}}_{\ell},\quad r^{\text{mlp}}_{\ell}\leftarrow r^{\text{mlp}}_{\ell}+\alpha\,v^{\text{mlp}}_{\ell}.\vspace{-5pt}$ (2)
@@
We set a maximum reasoning length of 8,192 tokens for GSM8K and 16,384 tokens for MATH-Level5. Upon reaching this limit, the model is prompted to finalize its answer immediately. We experiment on three reasoning models of varying sizes: deepseek-distill-{qwen-1.5B, llama3-8B, qwen-14B}.
@@
Figure 3 (Top) shows the effect of applying the attention-based direction $v_{\ell}^{\text{attn}}$ on GSM8K. We vary $\alpha$ from $-0.08$ (shorter reasoning) to $+0.08$ (longer reasoning). Across all models, increasing $\alpha$ extends the length of CoT (Figure 3, top right), indicating that $v_{\ell}^{\text{attn}}$ indeed encode reasoning-length attributes. In terms of accuracy, the larger 8B and 14B models improve when steered toward longer reasoning—particularly deepseek-distill-llama3-8B (orange line), which benefits most from positive steering with 10% accuracy improvement. In contrast, the smaller deepseek-distill-qwen-1.5B (blue line) model experiences a 10% drop in accuracy. Figure 3 (Bottom) presents the results for the more challenging MATH-Level5 dataset. Similar to GSM8K, our extracted directions effectively control reasoning length as expected, with negative $\alpha$ consistently leading to shorter CoT and positive $\alpha$ extending them. In terms of accuracy, shorter reasoning also consistently degrades performance. However, unlike GSM8K, there is no clear trend indicating that longer reasoning reliably enhances accuracy; while moderate positive $\alpha$ might provide some benefits, excessively long reasoning often negatively impacts performance. We also present results using the MLP-based direction $v_{\ell}^{\text{mlp}}$ in Appendix A.1, which exhibit similar trends.
@@ 4 ThinkEdit: Mitigate Overly Short Reasoning through Weight Editing
Building on the insights from Section 3.3, in this section, we propose ThinkEdit, an effective weight-editing method to mitigate overly short reasoning. We start by analyzing whether specific components within reasoning models significantly contribute to the phenomenon of short reasoning. Our focus is on pinpointing particular attention heads, as the attention mechanism plays a crucial role in information propagation across tokens. To explore this, we begin with an overview of the multi-head attention mechanism in Section 4.1, where we define the contribution of individual attention heads. Using this definition, we identify short reasoning heads in Section 4.2 and remove the short reasoning component from these heads in Section 4.3. Finally, in Section 4.4, we evaluate ThinkEdit and show that it effectively mitigates the overly short reasoning issue.
@@ 4.1 Overview of Attention-Head Structure
  • Q, K, and V Projections. Given a hidden-state ${r}\in\mathbb{R}^{T\times d}$ (for $T$ tokens), each head $h$ computes: $Q^{h}=rW_{q}^{h},\;K^{h}=rW_{k}^{h},\;V^{h}=rW_{v}^{h},$ where $Q^{h},K^{h},V^{h}\!\in\!\mathbb{R}^{T\times d_{h}}$. Each head $h$ has its own learnable projection matrices ${W}_{q}^{h},{W}_{k}^{h},{W}_{v}^{h}\!\in\!\mathbb{R}^{d\times d_{h}}$, which transform the hidden representation ${r}$ into query, key, and value vectors.
@@
    ${A}^{h}=\mathrm{softmax}\!\Bigl{(}\tfrac{{Q}^{h}({K}^{h})^{\top}}{\sqrt{d_{h}}}\Bigr{)}\,{V}^{h}\;\;\in\;\mathbb{R}^{T\times d_{h}}.\vspace{-5pt}$
@@
    $C^{h}\coloneq{A}^{h}{W}_{o}^{h}\;\in\;\mathbb{R}^{T\times d}.\vspace{-5pt}$ (3)
@@ 4.2 Identify Short Reasoning Attention Heads
For a response to problem $i$, let $\mathcal{T}_{i}$ be the set of token positions corresponding to the CoT, i.e., the tokens enclosed by <think> and </think> tags. Then, the overall average per-head contribution over all problems in the short reasoning dataset $\mathcal{D}_{\text{short}}$ is given by
$\overline{C}^{h}=\frac{1}{|\mathcal{D}_{\text{short}}|}\sum_{i\in\mathcal{D}_{\text{short}}}\left(\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}C^{h}(i,t)\right).\vspace{-5pt}$ (4)
Equation 4 first averages the per-head contributions $C^{h}(i,t)$ over the CoT token positions for each problem $i$ and then averages these values across all problems in $\mathcal{D}_{\text{short}}$. Recall that the reasoning length direction after an attention layer is defined as $v^{\text{attn}}_{\ell}=\overline{r}^{\text{attn}}_{\ell,\text{long}}-\overline{r}^{\text{attn}}_{\ell,\text{short}}$ in Equation 1. To quantify the short reasoning contribution of head $h$, we project $\overline{C}^{h}$ onto the negative of the reasoning length direction (i.e., the short reasoning direction). Using the unit vector $\hat{v}^{\text{attn}}_{\ell}=\frac{v^{\text{attn}}_{\ell}}{\|v^{\text{attn}}_{\ell}\|}$, we define the scalar projection as
$\overline{C}^{h}_{\textrm{short}}\;=\;\left\langle\overline{C}^{h},-\hat{v}^{\text{attn}}_{\ell}\right\rangle.\vspace{-5pt}$ (5)
@@ 4.3 Editing Short Reasoning Heads
We introduce how ThinkEdit effectively removes the short reasoning direction from the output projection matrices of the "short reasoning heads". Specifically, we identify the top 4% of attention heads with the largest $\overline{C}^{h}_{\text{short}}$ values (as defined in Section 4.2), marking them as short reasoning heads. Let $W_{o}^{h_{\ell}}\in\mathbb{R}^{d_{h}\times d}$ be the output projection matrix of head $h$ in layer $\ell$, and let $-\hat{v}^{\text{attn}}_{\ell}\in\mathbb{R}^{d}$ denote the short reasoning direction at layer $\ell$. We then update $W_{o}^{h_{\ell}}$ via:
$W_{o}^{h_{\ell}}\;\leftarrow\;W_{o}^{h_{\ell}}\Bigl{(}I\;-\;\bigl{(}-\hat{v}^{\text{attn}}_{\ell}\bigr{)}\bigl{(}-\hat{v}^{\text{attn}}_{\ell}\bigr{)}^{\!\top}\Bigr{)},\vspace{-5pt}$ (6)
where $I$ is the $d\times d$ identity matrix. Intuitively, this operation projects each row of $W_{o}^{h_{\ell}}$ onto the subspace orthogonal to $-\hat{v}^{\text{attn}}_{\ell}$, thereby removes the short reasoning component from the head’s output contribution. Unlike the approach in Section 3.3, which adds a fixed direction to activations regardless of the input, ThinkEdit modifies the weights of selected attention heads. This makes the adjustment input-dependent, allowing more precise control over reasoning length while preserving the model’s overall behavior.
@@ 4.4 Performance of Reasoning Models after ThinkEdit
We also explore different editing percentages and compare our approach to simply appending “Wait” to prompt longer reasoning; detailed results are provided in Appendix A.4. Additionally, results for deepseek-distill-qwen-32B are reported in Appendix A.5. Finally, We provide several concrete examples illustrating how ThinkEdit enhances reasoning quality in Appendix A.7.
@@ 5 Related Works
Controllable text generation has been explored across various domains Liang et al. (2024), with methods generally classified into training-time and inference-time control. These approaches aim to steer LLMs toward generating text with specific attributes while preserving fluency and coherence. Training-time control is achieved through fine-tuning Zeldes et al. (2020); Wei et al. (2022) or reinforcement learning Ouyang et al. (2022); Rafailov et al. (2023), leveraging diverse datasets to shape model behavior. Inference-time control encompasses prompt engineering Shin et al. (2020); Li and Liang (2021), representation engineering Subramani et al. (2022); Zou et al. (2023a); Konen et al. (2024), interpretable neuron intervention Sun et al. (2024b, c), and decoding-time interventions Dathathri et al. (2020), allowing flexible and efficient control without retraining.
@@
A growing body of research explores the role of attention heads and neurons within the Multi-Layer Perceptron (MLP) layers in shaping language model behavior. Studies such as Zhou et al. (2025); Zhao et al. (2025); Chen et al. (2024) examine how safety mechanisms are embedded in well-aligned models to defend against harmful prompts and jailbreak attacks Zou et al. (2023b); Liu et al. (2024); Sun et al. (2024a). Findings indicate that a small subset of attention heads and MLP neurons play a critical role in safety alignments. Similarly, research on hallucination mitigation has investigated the contributions of attention heads and MLP neurons. Ho et al. (2025) demonstrates that filtering out unreliable attention heads can reduce erroneous generations, while Yu et al. (2024) finds that activating subject-knowledge neurons helps maintain factual consistency.
@@ Acknowledgement
This research was supported by grants from NVIDIA and utilized NVIDIA’s A100 GPUs on Saturn Cloud. The authors are partially supported by Hellman Fellowship and Intel Rising Star Faculty Award.
@@ A.5 ThinkEdit Results for 32B Reasoning Model
A.6 Examples of Steering the Reasoning Length

### v4

@@ 1 Introduction
Recently, Reinforcement Learning (RL) has been applied to enhance Large Language Models (LLMs), equipping them with strong chain-of-thought (CoT) reasoning abilities Guo et al. (2025). These models, often referred to as reasoning models, first generate an intermediate reasoning process – a "thinking step" – where they reason step-by-step and then self-correct before producing a final response. As a result, they achieve remarkable improvement on mathematical reasoning tasks and demonstrate a strong ability to generate detailed CoT reasoning Jaech et al. (2024); Guo et al. (2025); Muennighoff et al. (2025).
@@ 3.2 Extracting Reasoning Length Directions
$\overline{r}^{x}_{\ell,y}=\frac{1}{|\mathcal{D}_{y}|}\sum_{i\in\mathcal{D}_{y}}\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}r^{x}_{\ell}(i,t),\vskip-5.0pt$
@@
These two vectors, $v^{\text{attn}}_{\ell},v^{\text{mlp}}_{\ell}\in\mathbb{R}^{d}$ (with $d$ denoting the hidden dimension), capture how the model’s representation differs when reasoning chains are notably longer or shorter. In the next section, we analyze how modifying these directions in the residual stream influences both reasoning length and overall model performance.
@@ 3.3 Effects of Reasoning-Length Direction
Let $\alpha$ be a scalar weight in the range [0.08, 0.08][-0.08,\,0.08]. For each layer $\ell$, we apply the following transformations:
$r^{\text{attn}}_{\ell}\leftarrow r^{\text{attn}}_{\ell}+\alpha\,v^{\text{attn}}_{\ell},\quad r^{\text{mlp}}_{\ell}\leftarrow r^{\text{mlp}}_{\ell}+\alpha\,v^{\text{mlp}}_{\ell}.\vskip-5.0pt$ (2)
@@
We set a maximum reasoning length of 8,192 tokens for GSM8K and 16,384 tokens for MATH-Level5. Upon reaching this limit, the model is prompted to finalize its answer immediately. We experiment on three reasoning models of varying sizes: deepseek-distill-qwen-1.5B, deepseek-distill-llama3-8B, and deepseek-
distill-qwen-14B
@@
Figure 3 (Top) shows the effect of applying the attention-based direction $v_{\ell}^{\text{attn}}$ on GSM8K. We vary $\alpha$ from 0.08-0.08 (shorter reasoning) to +0.08+0.08 (longer reasoning). Across all models, increasing $\alpha$ extends the length of CoT (Figure 3, top right), indicating that $v_{\ell}^{\text{attn}}$ indeed encode reasoning-length attributes. In terms of accuracy, the larger 8B and 14B models improve when steered toward longer reasoning—particularly deepseek-distill-llama3-8B (orange line), which benefits most from positive steering with 10% accuracy improvement. In contrast, the smaller deepseek-distill-qwen-1.5B (blue line) model experiences a 10% drop in accuracy. Figure 3 (Bottom) presents the results for the more challenging MATH-Level5 dataset. Similar to GSM8K, our extracted directions effectively control reasoning length as expected, with negative $\alpha$ consistently leading to shorter CoT and positive $\alpha$ extending them. In terms of accuracy, shorter reasoning also consistently degrades performance. However, unlike GSM8K, there is no clear trend indicating that longer reasoning reliably enhances accuracy; while moderate positive $\alpha$ might provide some benefits, excessively long reasoning often negatively impacts performance. We also present results using the MLP-based direction $v_{\ell}^{\text{mlp}}$ in Appendix A.1, which exhibit similar trends.
@@ 4 ThinkEdit: Mitigate Overly Short Reasoning through Weight Editing
Building on the insights from Section 3.3, in this section, we propose ThinkEdit, an effective weight-editing method to mitigate overly short reasoning. We start by analyzing whether specific components within reasoning models significantly contribute to the phenomenon of short reasoning. Our focus is on pinpointing particular attention heads, as the attention mechanism plays a crucial role in information propagation across tokens. To explore this, we begin with an overview of the multi-head attention mechanism in Section 4.1, where we define the contribution of individual attention heads. Using this definition, we identify short reasoning heads in Section 4.2 and remove the short reasoning component from these heads in Section 4.3. Finally, in Section 4.4, we evaluate ThinkEdit, and show that it effectively mitigates the overly short reasoning issue.
@@ 4.1 Overview of Attention-Head Structure
  • Q, K, and V Projections. Given a hidden-state ${r}\in\mathbb{R}^{T\times d}$ (for $T$ tokens), each head $h$ computes:
    $Q^{h}=rW_{q}^{h},\;K^{h}=rW_{k}^{h},\;V^{h}=rW_{v}^{h},$
    where $Q^{h},K^{h},V^{h}\!\in\!\mathbb{R}^{T\times d_{h}}$. Each head $h$ has its own learnable projection matrices ${W}_{q}^{h},{W}_{k}^{h},{W}_{v}^{h}\!\in\!\mathbb{R}^{d\times d_{h}}$, which transform the hidden representation r{r} into query, key, and value vectors.
@@
    ${A}^{h}=\mathrm{softmax}\!\Bigl(\tfrac{{Q}^{h}({K}^{h})^{\top}}{\sqrt{d_{h}}}\Bigr)\,{V}^{h}\;\;\in\;\mathbb{R}^{T\times d_{h}}.\vskip-5.0pt$
@@
    $C^{h}\coloneq{A}^{h}{W}_{o}^{h}\;\in\;\mathbb{R}^{T\times d}.\vskip-5.0pt$ (3)
@@ 4.2 Identify Short Reasoning Attention Heads
For a response to problem $i$, let $\mathcal{T}_{i}$ be the set of token positions corresponding to the CoT, i.e., the tokens enclosed by <think> and </think> tags. Then, the overall average per-head contribution over all problems in the short reasoning dataset $\mathcal{D}_{\text{short}}$ is given by $\overline{C}^{h}\in\mathbb{R}^{d}$:
$\overline{C}^{h}=\frac{1}{|\mathcal{D}_{\text{short}}|}\sum_{i\in\mathcal{D}_{\text{short}}}\left(\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}C^{h}(i,t)\right).\vskip-5.0pt$ (4)
Equation 4 first averages the per-head contributions $C^{h}(i,t)$ over the CoT token positions for each problem $i$ and then averages these values across all problems in $\mathcal{D}_{\text{short}}$. Recall that the reasoning length direction after an attention layer is defined as $v^{\text{attn}}_{\ell}=\overline{r}^{\text{attn}}_{\ell,\text{long}}-\overline{r}^{\text{attn}}_{\ell,\text{short}}$ in Equation 1, $v^{\text{attn}}_{\ell}\in\mathbb{R}^{d}$. To quantify the short reasoning contribution of head $h$, we project $\overline{C}^{h}$ onto the negative of the reasoning length direction (i.e., the short reasoning direction). Using the unit vector $\hat{v}^{\text{attn}}_{\ell}=\frac{v^{\text{attn}}_{\ell}}{\|v^{\text{attn}}_{\ell}\|}$, we define the scalar projection as $\overline{C}^{h}_{\textrm{short}}\in\mathbb{R}$:
$\overline{C}^{h}_{\textrm{short}}\;=\;\left\langle\overline{C}^{h},-\hat{v}^{\text{attn}}_{\ell}\right\rangle.\vskip-5.0pt$ (5)
@@ 4.3 Editing Short Reasoning Heads
We introduce how ThinkEdit effectively removes the short reasoning direction from the output projection matrices of the "short reasoning heads". Specifically, we identify the top 4% of attention heads with the largest $\overline{C}^{h}_{\text{short}}$ values (as defined in Section 4.2), marking them as short reasoning heads. Let $W_{o}^{h_{\ell}}\in\mathbb{R}^{d_{h}\times d}$ be the output projection matrix of head $h$ in layer $\ell$, and let $\hat{v}^{\text{attn}}_{\ell}\in\mathbb{R}^{d}$ denote the short reasoning direction at layer $\ell$. We then update $W_{o}^{h_{\ell}}$ via:
$W_{o}^{h_{\ell}}\;\leftarrow\;W_{o}^{h_{\ell}}\Bigl(I\;-\;\bigl(-\hat{v}^{\text{attn}}_{\ell}\bigr)\bigl(-\hat{v}^{\text{attn}}_{\ell}\bigr)^{\!\top}\Bigr),\vskip-5.0pt$ (6)
where $I$ is the $d\times d$ identity matrix. Intuitively, this operation projects each row of $W_{o}^{h_{\ell}}$ onto the subspace orthogonal to $\hat{v}^{\text{attn}}_{\ell}$, thereby removes the short reasoning component from the head’s output contribution. Unlike the approach in Section 3.3, which adds a fixed direction to activations regardless of the input, ThinkEdit modifies the weights of selected attention heads. This makes the adjustment input-dependent, allowing more precise control over reasoning length while preserving the model’s overall behavior.
@@ 4.4 Performance of Reasoning Models after ThinkEdit
In summary, ThinkEdit markedly improves model performance on short-reasoning instances and yields a substantial overall accuracy gain. We also explore different editing percentages and compare our approach to simply appending “Wait” to prompt longer reasoning; detailed results are provided in Appendix A.4. Additionally, results for deepseek-distill-qwen-32B are reported in Appendix A.5. To test the generality of our method, we further evaluate on non-math domains in Appendix A.6. Beyond demonstrating accuracy gains, Appendix A.7 examines how ThinkEdit shapes reasoning behaviors, revealing more explicit and structured chains of thought. Finally, We provide several concrete examples illustrating how ThinkEdit enhances reasoning quality in Appendix A.9.
@@ 5 Related Works
Controllable text generation has been explored across various domains Liang et al. (2024), with methods generally classified into training-time and inference-time control. These approaches aim to steer LLMs toward generating text with specific attributes while preserving fluency and coherence. Training-time control is achieved through fine-tuning Zeldes et al. (2020); Wei et al. (2022) or reinforcement learning Ouyang et al. (2022); Rafailov et al. (2023), leveraging diverse datasets to shape model behavior. Inference-time control encompasses prompt engineering Shin et al. (2020); Li and Liang (2021), representation engineering Subramani et al. (2022); Zou et al. (2023a); Konen et al. (2024); Oikarinen et al. (2025), interpretable neuron intervention through concept bottleneck models Sun et al. (2025b, 2024), and decoding-time interventions Dathathri et al. (2020), allowing flexible and efficient control without retraining.
@@
A growing body of research explores the role of attention heads and neurons within the Multi-Layer Perceptron (MLP) layers in shaping language model behavior. Studies such as Zhou et al. (2025); Zhao et al. (2025); Chen et al. (2024) examine how safety mechanisms are embedded in well-aligned models to defend against harmful prompts and jailbreak attacks Zou et al. (2023b); Liu et al. (2024); Sun et al. (2025a). Findings indicate that a small subset of attention heads and MLP neurons play a critical role in safety alignments. Similarly, research on hallucination mitigation has investigated the contributions of attention heads and MLP neurons. Ho et al. (2025) demonstrates that filtering out unreliable attention heads can reduce erroneous generations, while Yu et al. (2024) finds that activating subject-knowledge neurons helps maintain factual consistency. In Li et al. (2025), the authors design efficient and training-free machine skill unlearning techniques for LLMs through intervention and abstention.
@@ Acknowledgement
This research was supported by grants from NVIDIA and utilized NVIDIA’s A100 GPUs on Saturn Cloud. The authors are partially supported by Hellman Fellowship, Intel Rising Star Faculty Award, and National Science Foundation under Grant No. 2313105, 2430539.
@@ A.5 ThinkEdit Results for 32B Reasoning Model
A.6 ThinkEdit Results on Non-Math Domains
We include supplementary experiments on three non-math subjects from MMLU—High School Computer Science, Formal Logic, and Professional Accounting. These tasks do not directly require mathematical calculation, but they still demand structured reasoning and logical consistency. As shown in Tables 13 and 14, ThinkEdit not only improves overall accuracy but also mitigates failure cases that arise from overly short reasoning traces. This indicates that ThinkEdit can boost performance in reasoning-heavy domains beyond mathematics.
Model Variant MMLU HS CS MMLU Formal Logic MMLU Prof. Accounting
deepseek-qwen-32B Original 96.50 $\pm$ 1.18 93.49 $\pm$ 1.29 84.22 $\pm$ 1.05
ThinkEdit (4%) 97.20 $\pm$ 0.92 94.05 $\pm$ 1.80 83.40 $\pm$ 1.89
deepseek-qwen-14B Original 93.50 $\pm$ 1.58 91.27 $\pm$ 2.08 75.85 $\pm$ 1.89
ThinkEdit (4%) 94.80 $\pm$ 1.14 91.51 $\pm$ 1.98 77.09 $\pm$ 1.63
deepseek-llama3-8B Original 84.00 $\pm$ 3.83 63.41 $\pm$ 1.47 57.06 $\pm$ 1.48
ThinkEdit (4%) 88.90 $\pm$ 1.91 65.40 $\pm$ 2.85 57.52 $\pm$ 1.40
deepseek-qwen-1.5B Original 63.90 $\pm$ 4.38 51.27 $\pm$ 3.00 35.71 $\pm$ 2.85
ThinkEdit (4%) 68.30 $\pm$ 3.02 52.30 $\pm$ 3.07 37.09 $\pm$ 1.78
Table 13: Overall accuracy on MMLU non-math subjects.
Model Variant MMLU HS CS MMLU Formal Logic MMLU Prof. Accounting
deepseek-qwen-32B Original 98.00 / 98.00 / 97.50 85.00 / 88.33 / 91.60 89.29 / 87.86 / 87.86
ThinkEdit (4%) 100.00 / 99.00 / 99.50 88.33 / 88.33 / 92.00 92.86 / 91.07 / 91.96
deepseek-qwen-14B Original 96.00 / 94.00 / 96.00 78.33 / 85.83 / 90.40 82.14 / 82.50 / 84.82
ThinkEdit (4%) 100.00 / 99.00 / 99.50 85.00 / 90.00 / 92.00 90.00 / 91.43 / 90.36
deepseek-llama3-8B Original 72.00 / 76.00 / 81.00 75.00 / 75.83 / 74.00 70.71 / 70.71 / 67.32
ThinkEdit (4%) 96.00 / 95.00 / 96.00 80.00 / 77.50 / 81.20 75.00 / 74.64 / 70.54
deepseek-qwen-1.5B Original 66.00 / 66.00 / 71.00 48.33 / 51.67 / 61.60 32.86 / 32.14 / 33.21
ThinkEdit (4%) 92.00 / 89.00 / 84.00 60.00 / 61.67 / 67.20 43.57 / 43.21 / 40.36
Table 14: Accuracy on the top 5% / 10% / 20% shortest responses on MMLU non-math subjects.
A.7 Behavioral Analysis: ThinkEdit Encourages Deeper Reasoning
To better understand how ThinkEdit influences reasoning quality beyond final correctness, we analyze model behavior over all examples (correct and incorrect). We quantify (i) the average number of LaTeX equations written in the <think> trace, (ii) the percentage of examples whose final answer is explicitly boxed in <think>, and (iii) the average reasoning length (tokens) within <think>.
Across the four base models, ThinkEdit consistently produces more equations (+55.2% on average), boxes the final answer more often (+60.5%), and writes longer reasoning traces (+48.0%) compared to the original models. The effect is especially pronounced for smaller models (e.g., deepseek-qwen-1.5B: equations +83.9%, boxing +84.4%). These results indicate that the edits reliably encourage more explicit and structured chains of thought across the board.
Model Variant Equations in CoT Wrap Answer in CoT (%) Reasoning Length (tokens)
deepseek-qwen-32B Original 1.92 11.75 308.32
ThinkEdit (4%) 2.85 16.00 409.03
deepseek-qwen-14B Original 2.11 15.39 341.09
ThinkEdit (4%) 2.74 23.73 518.06
deepseek-llama3-8B Original 3.21 30.25 606.09
ThinkEdit (4%) 5.09 50.57 957.32
deepseek-qwen-1.5B Original 3.83 28.05 753.45
ThinkEdit (4%) 7.05 51.71 1125.64
Table 15: Behavioral comparison over all datasets. ThinkEdit generates more equations, is more likely to box the final answer, and produces longer reasoning traces than the original models, indicating more explicit and structured CoT across cases.
A.8 Examples of Steering the Reasoning Length